Given the differential equation y – 9y = - ett + 3e8t, y(0) = 0, y(0) = 4 Apply the Laplace Transform and solve for Y(8) =
Answer #1

y - 9y 8t at -e +3e on 2) S +3 SY(s)-4-9 YCS) - +4 (A+B+C+0) s +5/-14A-208-9C-89]+5 (21A +1238-90-90] y (0) = 0 , y(o)=4 AComponing co-efficient, A+B+C +D=0; 144 +206 +90 +80=-4 ; 214+.1238-90-90=-66 216A + (-216 B) +8167 72 D = 269 on solving, we